OpenStudy (anonymous):

.There are 200 questions on a 3 hr examination. Among these questions are 50 mathematics problems. It is suggested that twice as much time be spent on each Maths problem as for each other question. How many minutes should be spent on mathematics problems

OpenStudy (anonymous):

options A.36 B.72 C.60 D.100 Ans.B

OpenStudy (anonymous):

Generate equations in minutes preferably. total time is 180 minutes Let t be for normal and then 2t for math questions, So 2t*50+t*150=180

OpenStudy (anonymous):

superb ...:) ..thnkyou so mch @decidedlypretentious

OpenStudy (anonymous):

i am still not getting ans @decidedlypretentious

OpenStudy (anonymous):

got it, got it .:)

OpenStudy (anonymous):

100t + 150t =180 t=180/250

OpenStudy (anonymous):

i got ans, thnx:)

OpenStudy (anonymous):

thats for every quest. 300t is the answer